The purchased-equipment cost for a plant which produces pentaerythritol (solidfuelprocessing plant) is $300,000. The plant is to be an addition to an existing formaldehyde plant. 

The major part of the building cost will be for indoor construction, and the contractor’s fee will 

be 7 percent of the direct plant cost. All other costs are close to the average values found for 

typical chemical plants. On the basis of this information, estimate the following: (a) The total 

direct plant cost. (b) The fixed-capital investment. (c) The total capital investment.


1
Expert's answer
2021-05-12T15:59:29-0400

Total direct plant cost =$300, 000+fees


Contractor fees =0.07×300,000=21,000


i) Total direct plant cost =$321,000


TCI or FCI= Total direct plant cost × Ratio Factor


Using the Lang Factors for (Solid-Liquid Processing Plant)




ii) Fixed Capital Investment(FCI)


=321,000×4.3=$1,380,300


iii) Total Capital Investment (TCI)


=321,000 ×4.8=$1,540,800